OpenStudy (anonymous):

Find the solution to the differential equation, subject to the given initial condition.

OpenStudy (anonymous):

\[\frac{ dy }{ dx }+\frac{ y }{ 9 }=0, y(0)=16\]

myininaya (myininaya):

Did you try separation of variables.

OpenStudy (anonymous):

yeah i got dy+y/9=dx but im not really sure what to do after that

myininaya (myininaya):

I don't think you did the algebra there correctly.

myininaya (myininaya):

Why don't you try writing as f(x) dx=g(y) dy

myininaya (myininaya):

This is what I mean by separation of variables

OpenStudy (anonymous):

Ah. I have no idea what I'm doing then.

myininaya (myininaya):

Do you know how to undo the addition to isolate the dy/dx part?

myininaya (myininaya):

\[\frac{dy}{dx}+\frac{y}{9}=0\]

myininaya (myininaya):

You see addition you subtract.

myininaya (myininaya):

can you do that?

OpenStudy (anonymous):

so just subtract y/9 from both sides?

myininaya (myininaya):

That is how you get dy/dx by itself now you want g(y)dy=f(x)dx

OpenStudy (anonymous):

soo dy/y=-1/9 dx?

myininaya (myininaya):

that's right

myininaya (myininaya):

now integrate both sides

OpenStudy (anonymous):

okay, i'll try it. :)

OpenStudy (anonymous):

is ln(y)=-1/9x+c correct?

myininaya (myininaya):

now use the initial condition to find C you are given when x is 0 y is 16. Use that to find C.

OpenStudy (anonymous):

c=ln(16)

myininaya (myininaya):

yep

OpenStudy (anonymous):

dont I need to plug that back into the equation somewhere though?

myininaya (myininaya):

Yep C

OpenStudy (anonymous):

sooo ln(y)=-1/9x+ln(16)? i put that into the homework and it said it was wrong

myininaya (myininaya):

It might want you to solve for y. Also I would write it as ln|y|=-x/9+ln(16) I don't know if it would count off for not having those | | instead of ( ). But yeah if wants you to solve for y then solve for y.

OpenStudy (anonymous):

I don't even know how to go about doing that. ;_;

myininaya (myininaya):

So you don't know the inverse of natural log function?

OpenStudy (anonymous):

I don't remember it, no.

myininaya (myininaya):

what about the exponential function?

OpenStudy (anonymous):

uh wat.

myininaya (myininaya):

\[e^{\ln|x|}=x\]

myininaya (myininaya):

the exponential ( where base is e) function and natural log function are inverses

OpenStudy (anonymous):

ohhh yes, okay I remeber that now.

myininaya (myininaya):

use that to solve for y

myininaya (myininaya):

\[\text{ if } \ln|y|=f(x) \text{ then } y=e^{f(x)}\]

myininaya (myininaya):

steffw have you got that part?

myininaya (myininaya):

Once you do that. I will show you how to simplify the equation you get once you show me you have gotten that far.

OpenStudy (anonymous):

oops, i realized i was trying to solve for the wrong thing. xD i ended up with y=16e^-1/9x Thanks!

myininaya (myininaya):

omg you did good :)